OpenStudy (anonymous):

simplify. assume that no denominator equals 0. 54/72

OpenStudy (lgbasallote):

here's a hint \[\frac{54}{72} \implies \frac{54 \div 2}{72 \div 2}\]

OpenStudy (lgbasallote):

keep doing it until you cant do it no more

OpenStudy (anonymous):

3/4

OpenStudy (lgbasallote):

i think that's right...

OpenStudy (anonymous):

Yes.

OpenStudy (lgbasallote):

\[\frac{54 \div 18}{72 \div 18} = \frac 34\]

OpenStudy (anonymous):

did i do this right.. 10m/5m equal 2m

OpenStudy (lgbasallote):

no

OpenStudy (lgbasallote):

\[\frac{10m}{5m} \implies \frac{10 \cancel m}{5 \cancel m}\]

OpenStudy (anonymous):

so it would just be 2

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

okay thanks

OpenStudy (lgbasallote):

welcome